A two digit number is 4 times the sum of its digits and twice the product of its digit.FIND THE NUMBER
Answer by ankor@dixie-net.com(22740)

two digits, x & y
10x + y = the number
:
A two digit number is 4 times the sum of its digits
10x + y = 4(x+y)
10x + y = 4x + 4y
10x - 4x = 4y - y
6x = 3y
divide both sides by 3
2x = y
" and twice the product of its digit."
10x + y = 2xy
Replace y with 2x
10x + 2x = 2*x*2x
12x = 4x^2
Divide both sides by 4x
3 = x
then
y = 2(3)
y = 6
:
the number = 36
:
You can confirm this in both statement:
36 = 4(3+6)
and
36 = 2(3*6)